# Distributed Multisensor ISAC

### v1

@@ II-B Uplink / Downlink Sensing
Compared to infrastructure-only sensing, uplink / downlink sensing utilizes an UE directly as an illuminator or a sensor. The respective complementary sensor or illuminator side is taken by the BS, see Figure 3. Now, the sensing result depends on the availability, position and dynamics of the UEs, e.g. cars, drones or sniffing sensor nodes that are just deployed on demand or moving. In any case, the sensors and the illuminator are not directly connected by a fixed link as we have postulated in cases of the distributed infrastructure-only sensing. However, we assume the UE is booked into the Uplink (UL) or DL of the base station. Therefore, the sensing geometry is bistatic in general and corresponds to the “natural” arrangement of UE vs. infrastructure in mobile communication. This situation has some striking similarities to the passive radar principle, which is also called Passive Coherent Location (PCL) and well known since decades [ref29]. PCL relies on a so called transmitter of opportunity, e.g. a digital broadcasting DVB-T or analog FM transmitter. The PCL sensor tries to receive the clean transmit signal which is used as a correlation reference for the surveillance signal. Therefore, a conventional PCL sensor typically has two receiver channels, a highly directive channel delivering the multipath-free copy of the Tx signal and an omnidirectional surveillance channel containing the target response. Besides of the target related diversity gain of the bistatic sensing geometry, PCL offers the advantages that Tx  - Rx synchronization is no longer an issue as we only need to know the ToF difference between the reference and the surveillance channel at the receiver, which we call Excess Time of Flight (eToF). Of course, we have to make sure that the transmit reference is received via LOS and we need to know the position of the (mobile) UE.
@@
In continuation of the PCL concept we have proposed CPCL in [ref1]. The discussion about CPCL will run through the entire paper. It turns out that the advantages of CPCL are striking and manifold. CPCL is the way to fully integrate sensing functionality into the radio access scheme for communication. It allows to use the pilot and reference signals for Tx  / Rx frame synchronization and cooperation based on feedback signaling. It also paves the way to utilize the full communications data payload for radar sensing, while taking full advantage of the flexibility of 5G New Radio (NR) access schemes for signal adaption.
@@
A mobile ISAC sensor in the DL can already take great benefits even if the sensor is only passive (sniffing) and does not act as a full authorized communication user. In this reduced CPCL-mode, a frame synchronized receiver can take full advantage of Fast Fourier Transform (FFT) based frequency domain system identification methods [ref30, ref96] over classical PCL. This includes cyclic prefix removal and lower variance since intercarrier leakage almost disappears. Moreover, the received signal can be equalized which supports transmit reference signal recovery without requiring a dedicated (authorized) receiver channel. In the sniffer mode, the DL plays the role as broadcast sensing channel even if it is not a broadcast channel for communication. However, as there is no active interaction with the illuminating node, the network resources cannot be used for data fusion. If the mobile ISAC sensor is a full authorized UE, the communication functions on the network level can be used for data transfer, data fusion (with access to computational resources), predistortion at transmitter side and resource allocation which makes CPCL most attractive if mobile or deployable UE devices are included.
@@
Actually, there are considerable differences between UL and DL ISAC in terms of multisensor access and reasonable application scenarios. While a mobile sensor, being a fully connected UE or just a sniffing node, can use the complete multiuser communication payload transmitted by the gNB in the DL for broadcast ISAC illumination, the situation in the UL is different. Access in UL has to be designed as orthogonal MS ISAC since the target illumination is facilitated by the multiple transmitting UEs which are at different positions. However, this immediately gives the gNB multiple concurrent measurements, hence multiple ellipses, for target localization (assuming that the target is in the sensing coverage area of both illuminating UEs). But what about localization in the DL case? From where do we get the multiple results? There are two options: In case of multiple UEs (respectively mobile sniffing sensors) receiving sensing signals, they would have to share their sensing data for data fusion, which would need dedicated communication channels. The other option is that the mobile sensing node should look out for another illuminating gNB. If this auxiliary network node is used for CPCL in its reduced form, just as a sniffer node without communication connectivity to the illuminator, it does not need to decode data and no communication resources are consumed. However, the respective illumination signal has to be taken as it is and no cooperation with the transmitter (e.g., no adaptive modification of the transmit signal) is possible. Required metadata can be transferred by the primary link. Obviously, the auxiliary illumination source can be different in its radio parameters from the primary ISAC link, e.g. located at a different frequency or with different bandwidth and modulation. This would open interesting perspectives for radar sensing. Further drawbacks of UL sensing result from the lower transmit power of the mobile UE and from lower duty cycle in Time Division Duplex (TDD) for target illumination.
@@
ISAC in down link has the potential to complement the already very well established auto motive radar and LiDAR sensors in cars. We do not assume that ISAC can replace these sensors, but ISAC can support the “ego view” that only emanates from the illuminators and sensors installed on the cars. This makes the benefit immediately visible for the road user. In DL ISAC access, the additional illumination by infrastructure nodes can help to better recognize the local environment of the car. The bistatic radar architecture already provides an advantage with its inherent diversity advantage that supports an extended field of view and more comprehensive perception. Of course, the biggest gain will arise if all sensor data available in the car, including ego radar and LiDAR, will be fused. ISAC can provide additional road safety as it is an independent system. This can be an advantage if the density of automotive radars undergoes some saturation because of mutual interference (as automotive radar does not include an orthogonal multisensor access scheme). Other areas of application can be comprehensive perception, navigation and collision avoidance for Automated Guided Vehicels (AGV) in warehouses, extended manufacturing facilities, large logistic facilities, reloading sides, etc. As the mobile sensors can rely on the more powerful base station to illuminate the scene, promising applications are conceivable in which the mobile sensor explores its surroundings. There is a similar advantage in detecting low-flying drones that might be hidden behind buildings or other terrain structures. They become visible when illuminated by base stations on the ground and observed by high-flying sensing drones. Also in this scenario, the high transmission power and ubiquitous availability of base stations allow better illumination of the scenario, while the sensing drone only needs to carry a receiver when observing the scenario.
The UL scenario seems to be more appropriate to support the operator respectively regulatory view described above for infrastructure only sensing. The mobile, distributed illuminating node can help to get a better illuminated scenario. In special critical situations, the mobile node can be assigned a specific illumination mission.
Figure 3: MS ISAC with mobile UE included in the downlink or uplink.
@@ II-C Sidelink Based Sensing in Mesh Networks
Using the SL for integrated sensing in dense traffic scenarios can substantially extend the situational awareness compared to the established ego-view oriented automotive radars. This gain becomes also available if no infrastructure is in reach. It results from the spatial diversity of the meshed, distributed MIMO radar network, which is spanned by the cluster of neighboring vehicles. Compared to automotive radar, the lower frequencies (FR1) offer wider coverage due to lower path loss and less shadowing, and, to some extent, enables the visibility of targets in non-LOS regions. Hence, meshed SL ISAC has the potential to expand the ego-view. It will be an important step towards cooperative sensing. An advantage over the pure exchange of ego data can be that ISAC data result from using a common, standardized radio interface. Therefore, car manufactures maybe less reluctant to contribute to this type of a cooperation since they do not need to share their proprietary ego-sensing data. ISAC radar sensing in combination with SL communication can also be used to increase trustworthiness for communication since it can help to identify fake transmitters. For instance, a malign SL node trying to fool the other connected road users by sending wrong Cooperative Awareness Messages (CAM), can be identified by comparing the pretended CAM information with the radar ground truth measured by ISAC (relative distance or speed).
Besides of application of meshed ISAC sensing using the SL in public road traffic, also other applications with mobile clouds of connected vehicles are on the horizon. Examples are autonomously acting AGV s or swarms of drones. Application of drones for real-time monitoring of ground activities, for or the purpose of law enforcement, search and rescue, technical inspections, environmental monitoring, etc. is already widespread. For instance, drones can help to increase safety of big public events by early recognizing mass panic and recognizing critical streams of movement, detect blocked or closed exits or even guide people to open exit routes. Auxiliary ISAC sensing can help to extend the perceptive capabilities by the radio domain. Radar sensing can help when optical sensing fails because of bad whether condition or missing natural optical illumination. With radar sensing we may easier get an overview picture and measure relative velocities. If, in addition to radar detection, ISAC also detects intentional or unintentional radio emissions from the ground, we can get a much more comprehensive picture of the situation on the ground. Moreover, besides sensing of ground activities, SL based sensing can help to localize relative positions of drones with the swarm detect obstacles for collision avoidance (see Figure 4).
@@ III-A Distributed MS MIMO ISAC
The three basic architectures of MS ISAC systems and the corresponding use cases discussed in the previous chapter have already demonstrated the advantages of distributed MIMO ISAC concepts. This already included several bistatic, i.e. multistatic, sensing links. It also became clear that the multisensor radio access procedures, the allocation and scheduling of radio resources and synchronization issues will be very diverse and extensive. Ultimately, we need a comprehensive architecture for accessing and coordinating multiple sensors. The purpose of this chapter is to outline a generic MS MIMO ISAC architecture and to highlight the potential basic system design. In terms of radar community, we obtain a distributed MIMO radar architecture [ref10], [ref10b], which differs well from a MIMO radar with collocated antennas [ref10a]. The latter is a favored solution for quasi-monostatic radio access, as it combines the advantage of separated Tx and Rx antennas arrays on the same platform for reduced Tx to Rx interference with a reduced hardware overhead for forming a big virtual array from smaller, respectively sparse Tx/Rx subarrays at the expense of some transmit multiplexing scheme. For a more comprehensive discussion of collocated MIMO radar readers are referred to [ref10b] and the references given there. However, distributed MIMO radar corresponds much better to MS MIMO ISAC as discussed already in the section above. Another striking equivalence becomes obvious when MS-MIMO ISAC is compared to MU MIMO [ref31], which was a big step forward in mobile communications compared to single-user MIMO by adding multiuser capability in the wireless realm. Therefore, it seems worthwhile to consider advanced MU access schemes in mobile communication for application to MS ISAC.
@@
From Figure 5 it becomes clear, that collecting the full distributed NxN MIMO matrix with the $N$-node MS-ISAC network requires $N$ monostatic and $N^{2}-N$ bistatic measurements. Each measurement consists of a normalized cross-correlation function of the respective received signal $y_{n}(t)$ with the transmitted excitation signal $x_{m}(t)$ (the correlation reference) between nodes $m$ and $n$, which is regarded as an estimate of the corresponding Tx-to-Rx Channel Impulse Response (CIR) $\bm{h}_{m,n}(t)$, respectively Frequency Response Functions (FRF) $\bm{H}_{m,n}(f)$. Equation 1 depicts the full distributed I/O MIMO matrix according to fig. 5. The elements in the main diagonal result from the monostatic measurements whereas the entries in the upper and the lower triangular part corresponds to bistatic measurements in the Forward (FW) and Backward (BW) link (UL and DL, respectively). If FW and BW connections are reciprocal, one half of the bistatic part can be regarded as redundant. However, since radio nodes can be heterogeneous in the FW, BW and access mode (hence non-reciprocal) we can get up to $N^{2}$ independent measurements. This already clearly shows the difference of multistatic vs. multiple monostatic measurements. Why do we need so many measurements? The advantages are the resulting number of spatial degrees of freedom and diversity. While the former is well defined by the sensor-target geometry, the latter has a statistical meaning. This means that we can select the measurements with the highest geometric relevance for localization and omit the others (e.g., the monostatic links if there is a lack of full duplex radio interface), or we can use the seemingly redundant measurements for further location variance and bias reduction measures and better target recognition. Surplus measurements, e.g., can also be used to detect targets in cases of missing LOS, respectively target shadowing. Another advantage of distributed MS ISAC comes from increased spatial coverage and robustness since the sensor nodes act independently and can complement and substitute each other. It seems clear that access coordination between distributed links and collaboration may become more important for MS-MIMO ISAC than for MU MIMO communication, as target localization and tracking is bound to a clear multidimensional geometry and subjected to real-time conditions.
@@
where $X_{n}$ denotes the transmit signal of node $n$ and $Y_{m}$ the receive signal at node $m$. Hence, (1) essentially denotes the estimated FRFs between nodes $m$ and $n$, corresponding to Figure 5.
@@ III-B Heterogenous Radio Nodes
The radio nodes of the distributed MS-MIMO ISAC network in Figure 5. can be heterogeneous in many aspects. Heterogeneity can occur on the level of radio parameters or on the level of network parameters. In terms of the radio parameters, we may see differences w.r.t. frequency range, bandwidth, transmit and receive access schemes, antennas, polarimetric access, etc. These parameters are related to wave propagation aspects and to design issues of the radio interface. Since mobile radio frequencies range from well below $1\text{\,}\mathrm{GHz}\text{/}$ up to millimeter wave frequencies (perhaps more than $100\text{\,}\mathrm{GHz}\text{/}$ in the future), we can take advantage of the different scattering, diffraction and penetration characteristics in this wide span of frequencies. Roughly speaking, the higher the frequency, the more optical-like the propagation mechanisms become. At lower frequencies, we observe more diffraction and diffuse scattering, also materials penetration becomes possible. At higher frequencies, objects are more likely to act as reflectors, although surface roughness may become more important and obstructing objects may cause more blockage. Moreover, polarization selectivity of target details may become stronger. Finally, the same relative speed results in respectively higher Doppler shift. As far as the radio interface is concerned, the antennas play the most important role. Since the directivity depends to a certain extent on the size of the antenna in relation to the wavelength, the antennas tend to be smaller for the same directivity (or gain). However, link attenuation increases because of the smaller effective antenna area. This leads to the use of larger and therefore more directional antennas, which must be controllable in mobile applications.
@@
From a technical point of view, the combination of different frequencies offers many advantages, but also poses challenges. While only smaller frequency distances can still be handled by a single Up/Down Converter (UDC) radio interfaces, carrier aggregation and wider frequency spacing (e.g. FR1 and FR2) may require multiple RF chains or non-heterogeneous radio nodes. While widely separated frequency blocks can increase delay resolution a lot, coherent processing would be necessary which is possible if the frequency blocks belong to the same UDC channel (interband processing). If the frequency blocks belong to different UDC channels, non-coherent combining of distant frequency bands is easier to achieve and can still help to increase target resolution, may make detection more robust, and reveal different features for target recognition by frequency diversity. Another advantage frequency diversity results from the interplay between coverage and resolution. Lower frequencies enable wider coverage with less directional antennas. Higher frequencies with its wider bandwidth channels support higher delay (resp. range) resolution. At the same time, beamforming becomes possible which allows spatial filtering (to suppress clutter) and increased SNR. However, any beamforming also requires beam search, resp. scanning procedures, which may be seen as too tedious and cumbersome for fast target acquisition. Obviously, a cooperation of low frequency nodes for target acquisition and high frequency nodes for precise localization, recognition and tracking seems to be an advantage. This may include beam search instruction at higher frequencies from measurements at lower frequencies.
@@
Concerning the antennas and antenna arrays used, a wide variety of differences can occur. Here we will only mention some. The difference between linear/planar vs. circular/cylindrical array, e.g., decides about full azimuth angle coverage which seems be more important for sensing than for communications. Moreover, elevation coverage matters if targets are expected in the air (e.g. drones). Important is also the difference between analog (phased array) and digital beamforming. While the former is capable only for beamforming, the latter has the potential of direction estimation. Reason is that estimation of direction is based on mutual correlation processing of array outputs. Analog beamforming (which seems dominating for higher frequencies) embodies a linear weight and sum processor that does not allow access to the individual antenna outputs for correlation processing. Full polarimetric processing requires polarization selective antennas and two corresponding UDC channels and Tx and Rx side.
@@
The sniffer nodes discussed above may fall out of the ordinary even more. At first, a sniffer will act as receive only. But even if it is hosted by a gNB, it can cover extended frequency bands outside the host’s assigned bands. Receiving off-system signals would help to collect additional ISAC data in the reduced CPCL mode (like a synchronized passive radar) relative to other coexisting mobile radio networks that produce valuable illuminating signal (could also include DVB-T broadcast). At the same time, the sniffer can be used to detect and localize alien emitters. This feature opens up other wide fields of ISAC applications and can be an interesting add-on for some use cases.
@@
If the sniffer is integrated into a distributed base station architecture by using a fixed wireless, RFoF, or optical fronthaul link, it can be directly incorporated into the internal gNB base-band processing regime and enjoy all advantages of CPCL processing. However, if it is equipped with its own subscriber identity (SIM card), it can act as normal UE and can be deployed more easily at preferred locations. Now it can interact with the hosting gNB via the normal UE UL/DL radio interface. However, the CPCL advantages in terms of carrier and clock reference synchronization still apply.

### v2

@@ II-B Uplink / Downlink Sensing
Compared to infrastructure-only sensing, uplink / downlink sensing utilizes a UE directly as an illuminator or a sensor in the sensing cycle, where the BS acts as complementary sensor or illuminator, see Figure 3. Now, the sensing result depends on the availability, position and dynamics of the UEs, e.g. cars, drones or sniffing sensor nodes that are deployed only on demand or moving. In any case, the sensors and the illuminator are not directly connected by a fixed link as we have postulated in cases of the distributed infrastructure-only sensing. However, we assume the UE is booked into the Uplink (UL) or DL of the base station. Therefore, the sensing geometry is bistatic in general and corresponds to the “natural” arrangement of UE vs. infrastructure in mobile communication. This situation has some striking similarities to the passive radar principle, which is also called Passive Coherent Location (PCL) and well known for decades [ref29]. PCL relies on a so called transmitter of opportunity, e.g. a digital broadcasting DVB-T or analog FM transmitter. The PCL sensor tries to receive the clean transmit signal which is used as a correlation reference for the surveillance signal. Therefore, a conventional PCL sensor typically has two receiver channels, a highly directive channel delivering the multipath-free copy of the Tx signal and an omnidirectional surveillance channel containing the target response. Besides of the target related diversity gain of the bistatic sensing geometry, PCL offers the advantage that Tx  - Rx synchronization is no longer an issue as we only need to know the ToF difference between the reference and the surveillance channel at the receiver, which we call eToF. Of course, we have to make sure that the transmit reference is received via LOS and we need to know the position of both, the BS and the (mobile) UE.
@@
In continuation of the PCL concept we have proposed CPCL in [ref1]. The discussion about CPCL will run through the entire paper. It turns out that the advantages of CPCL are striking and manifold. CPCL is a way to fully integrate sensing functionality into the radio access scheme for communication. It allows to use the pilot and reference signals for Tx  / Rx frame synchronization and cooperation based on feedback signaling. It also paves the way to utilize the complete communications data payload for radar sensing, while taking full advantage of the flexibility of 5G New Radio (NR) access schemes for signal adaption.
@@
A mobile ISAC sensor in the DL can already obtain benefits even if the sensor is only passively sniffing and does not act as a fully authorized communication user. In this reduced CPCL-mode, a frame synchronized receiver can take full advantage of Fast Fourier Transform (FFT) based frequency domain system identification methods [ref30, ref96] over classical PCL. This includes cyclic prefix removal and thus lower variance since intercarrier leakage almost disappears. Moreover, the received signal can be equalized which supports transmit reference signal recovery without requiring a dedicated (authorized) receiver channel. In the sniffer mode, the DL plays the role as broadcast sensing channel even if it is not a broadcast channel for communication. However, as there is no active interaction with the illuminating node, the network resources cannot be used for data fusion. If the mobile ISAC sensor is a fully authorized UE, the communication functions on the network level can be used for reporting of sensing results and subsequent data fusion (with access to computational resources), predistortion at transmitter side and resource allocation which makes CPCL most attractive if mobile or deployable UE devices are included.
@@
Actually, there are considerable differences between UL and DL ISAC in terms of multisensor access and reasonable application scenarios. While a mobile sensor, being a fully connected UE or just a sniffing node, can use the complete multiuser communication payload transmitted by the gNB in the DL for broadcast ISAC illumination, the situation in the UL is different. Access in UL has to be designed orthogonally since the target illumination is facilitated by the multiple transmitting UEs which are at different positions. However, this immediately gives the gNB multiple concurrent measurements, hence multiple ellipses, for target localization (assuming that the target is in the sensing coverage area of both illuminating UEs). But what about localization in the DL case? From where do we get the multiple results? There are two options: In case of multiple UEs (respectively mobile sniffing sensors) receiving sensing signals, they would have to share their sensing data for data fusion, which would need dedicated communication channels. The other option is that the mobile sensing node should look out for another illuminating gNB. If this auxiliary network node is used for CPCL in its reduced form, just as a sniffer node without communication connectivity to the illuminator, it does not need to decode data and no communication resources are consumed. However, the respective illumination signal has to be taken as it is and no cooperation with the transmitter (e.g., no adaptive modification of the transmit signal) is possible. Required metadata can be transferred by the primary link. Obviously, the auxiliary illumination source can be different in its radio parameters from the primary ISAC link, e.g. located at a different frequency or with different bandwidth and modulation. This would open interesting perspectives for radar sensing. Further drawbacks of UL sensing result from the lower transmit power of the mobile UE and from lower duty cycle in Time Division Duplex (TDD) for target illumination.
@@
ISAC in down link has the potential to complement the already very well established auto motive radar and LiDAR sensors in cars. We do not assume that ISAC can replace these sensors, but ISAC can support the “ego view” that only emanates from the illuminators and sensors installed on the cars. This renders the benefit immediately visible for the road user. In DL ISAC access, the additional illumination by infrastructure nodes can help to better recognize the local environment of the car. The bistatic radar architecture already provides an advantage with its inherent diversity advantage that supports an extended field of view and more comprehensive perception. Of course, the biggest gain will arise if all sensor data available in the car, including ego radar and LiDAR, can be fused. ISAC can provide additional road safety as it is an independent system, which can be an advantage if the density of automotive radars undergoes some saturation because of mutual interference as an automotive radar does not include an orthogonal multisensor access scheme. Other areas of application can be comprehensive perception, navigation and collision avoidance for Automated Guided Vehicels (AGV) in warehouses, extended manufacturing facilities, large logistic facilities, reloading sides, etc. As the mobile sensors can rely on the more powerful base station to illuminate the scene, promising applications are conceivable in which the mobile sensor explores its surroundings. There is a similar advantage in detecting low-flying drones that might be hidden behind buildings or other terrain structures. They become visible when illuminated by base stations on the ground and observed by high-flying sensing drones. Also in this scenario, the high transmission power and ubiquitous availability of base stations allow better illumination of the scenario, while the sensing drone only needs to carry a receiver when observing the scenario.
The UL scenario seems to be more appropriate to support the operator respectively regulatory view described above for infrastructure only sensing. The mobile, distributed node can help to get a better illuminated scenario. In special critical situations, the mobile node can be assigned a specific illumination/sensing mission.
Figure 3: MS-ISAC with mobile UE included in the downlink or uplink.
@@ II-C Sidelink Based Sensing in Mesh Networks
Using the SL for integrated sensing in dense traffic scenarios can substantially extend the situational awareness compared to the established ego-view oriented automotive radars even if no infrastructure is in reach. It results from the spatial diversity of the meshed, distributed MIMO radar network, which is spanned by the cluster of neighboring vehicles. Compared to automotive radar, the lower frequencies (FR1) offer wider coverage due to lower path loss and less shadowing, and, to some extent, enables the visibility of targets in non-LOS regions. Hence, meshed SL ISAC has the potential to expand the ego-view. It will be a step towards cooperative sensing. An advantage over the pure exchange of ego data can be that ISAC data result from using a common, standardized radio interface. Therefore, car manufactures maybe less reluctant to contribute to this type of a cooperation since they do not need to share their proprietary ego-sensing data. ISAC radar sensing in combination with SL communication can also be used to increase trustworthiness for communication since it can help to identify fake transmitters. For instance, a malign SL node trying to fool the other connected road users by sending wrong Cooperative Awareness Messages (CAM), can be identified by comparing the pretended CAM information with the radar ground truth measured by ISAC (relative distance or speed).
Besides meshed ISAC sensing using the SL in public road traffic, other applications with mobile clouds of connected vehicles are on the horizon. Examples are autonomously acting AGVs or swarms of drones. Application of drones for real-time monitoring of ground activities, for or the purpose of law enforcement, search and rescue, technical inspections, environmental monitoring, etc. is already widespread. For instance, drones can help to increase safety of big public events by early recognizing mass panics and critical streams of movement, detecting blocked or closed exits or even guiding people to open exit routes. Auxiliary ISAC sensing can help to extend the perceptive capabilities by the radio domain. Radar sensing can help when optical sensing fails because of bad whether condition or missing natural optical illumination. With radar sensing we may get an overview picture more easily and measure relative velocities. If, in addition to radar, ISAC also detects intentional or unintentional radio emissions from the ground, we can get a more comprehensive picture of the situation on the ground. Moreover, besides sensing of ground activities, SL based sensing can help to localize relative positions of drones towards the swarm for collision avoidance (see Figure 4).
@@ III-A Distributed MS MIMO ISAC
The three basic architectures of MS-ISAC systems and the corresponding use cases discussed in the previous chapter have already demonstrated the advantages of distributed MIMO ISAC concepts. This already included several bistatic, i.e. multistatic, sensing links. It also became clear that the multisensor radio access procedures, the allocation and scheduling of radio resources and synchronization issues will be diverse and extensive. Ultimately, we need a comprehensive architecture for accessing and coordinating multiple sensors. The purpose of this chapter is to outline a generic MS MIMO ISAC architecture and to highlight the potential basic system design. In terms of the radar community, we obtain a distributed MIMO radar architecture [ref10, ref10b], which differs well from a MIMO radar with collocated antennas [ref10a]. The latter is a favored solution for quasi-monostatic radio access. It combines the advantage of separated Tx and Rx antennas arrays on the same platform for reduced Tx to Rx interference with a reduced hardware overhead for forming a big virtual array from smaller, respectively sparse Tx/Rx subarrays at the expense of some transmit multiplexing scheme. For a more comprehensive discussion of collocated MIMO radar readers are referred to [ref10b] and the references given therein. However, distributed MIMO radar corresponds much better to MS MIMO ISAC as discussed already in the section above. Another striking equivalence becomes obvious when MS-MIMO ISAC is compared to MU MIMO [ref31], which was a big step forward in mobile communications compared to single-user MIMO by adding multiuser capability in the wireless realm. Therefore, it seems worthwhile to consider advanced MU access schemes in mobile communication for application to MS-ISAC.
@@
From Figure 5 it becomes clear that collecting the full distributed $N\times N$ MIMO matrix with the $N$-node MS-ISAC network requires $N$ monostatic and $N^{2}-N$ bistatic measurements. Each measurement consists of a normalized cross-correlation function of the respective received signal $y_{n}(t)$ with the transmitted excitation signal $x_{m}(t)$ (the correlation reference) between nodes $m$ and $n$, which is regarded as an estimate of the corresponding Tx-to-Rx Channel Impulse Response (CIR) $\bm{h}_{m,n}(t)$, respectively Frequency Response Functions (FRF) $\bm{H}_{m,n}(f)$. Equation 1 depicts the full distributed I/O MIMO matrix according to fig. 5. The elements in the main diagonal result from the monostatic measurements whereas the entries in the upper and the lower triangular part corresponds to bistatic measurements in the Forward (FW) and Backward (BW) link (UL and DL, respectively). If FW and BW connections are reciprocal, one half of the bistatic part can be regarded as redundant. However, since radio nodes can be heterogeneous in the FW, BW and access mode (hence non-reciprocal) we can get up to $N^{2}$ independent measurements. This already shows the difference of multistatic vs. multiple monostatic measurements. Why do we need so many measurements? The advantages are the resulting number of spatial degrees of freedom and diversity. While the former is well defined by the sensor-target geometry, the latter has a statistical meaning. We can select the measurements with the highest geometric relevance for localization and omit the others (e.g., the monostatic links if there is a lack of full duplex radio interface), or we can use the seemingly redundant measurements for further location variance and bias reduction measures and improve target recognition. Surplus measurements can also be used to detect targets in cases of missing LOS, respectively target shadowing. Another advantage of distributed MS-ISAC comes from increased spatial coverage and robustness since the sensor nodes act independently and can complement and substitute each other. Access coordination between distributed links and collaboration may become more important for MS-MIMO ISAC than for MU MIMO communication, as target localization and tracking is bound to a clear multidimensional geometry and subjected to real-time conditions.
@@
and $X_{n}$ denotes the transmit signal of node $n$ and $Y_{m}$ the receive signal at node $m$. Hence, (1) essentially denotes the estimated FRFs between nodes $m$ and $n$, corresponding to Figure 5.
@@ III-B Heterogenous Radio Nodes
The radio nodes of the distributed MS-MIMO ISAC network in Figure 5 can be heterogeneous in many aspects. Heterogeneity can occur on the level of radio parameters or on the level of network parameters. In terms of the radio parameters, we may see differences w.r.t. frequency range, bandwidth, transmit and receive access schemes, antennas, polarimetric access, etc. These parameters are related to wave propagation aspects and to design issues of the radio interface. Since mobile radio frequencies range from well below $1\text{\,}\mathrm{GHz}\text{/}$ up to millimeter wave frequencies (perhaps more than $100\text{\,}\mathrm{GHz}\text{/}$ in the future), we can take advantage of the different scattering, diffraction and penetration characteristics in this wide span of frequencies. Roughly speaking, the higher the frequency, the more optical-like the propagation mechanisms become. At lower frequencies, we observe more diffraction and diffuse scattering, also material penetration becomes possible. At higher frequencies, objects are more likely to act as reflectors, although surface roughness may become more important and obstructing objects may cause more blockage, and polarization selectivity of target details may become stronger. Finally, the same relative speed results in respectively higher Doppler shift. As far as the radio interface is concerned, the antennas play the most important role. Since the directivity depends to a certain extent on the size of the antenna in relation to the wavelength, the antennas tend to be smaller for the same directivity (or gain). However, link attenuation increases because of the smaller effective antenna area. This leads to the use of larger and therefore more directional antennas, which must be controllable in mobile applications.
@@
From a technical point of view, the combination of different frequencies offers many advantages, but also poses challenges. While only smaller frequency distances can still be handled by a single Up/Down Converter (UDC) radio interface, carrier aggregation and wider frequency spacing (e.g. FR1 and FR2) may require multiple RF chains or non-heterogeneous radio nodes. While widely separated frequency blocks can increase delay resolution a lot, coherent processing would be necessary which is possible if the frequency blocks belong to the same UDC channel (interband processing). If the frequency blocks belong to different UDC channels, non-coherent combining of distant frequency bands is easier to achieve and can still help to increase target resolution, may make detection more robust, and reveal different features for target recognition by frequency diversity. Another advantage of frequency diversity results from the interplay between coverage and resolution. Lower frequencies enable wider coverage with less directional antennas. Higher frequencies with their wider bandwidth channels support higher delay (resp. range) resolution. At the same time, beamforming becomes possible which allows spatial filtering (to suppress clutter) and increase SNR. However, any beamforming also requires beam search, resp. scanning procedures, which may be seen as too tedious and cumbersome for fast target acquisition. Obviously, a cooperation of low frequency nodes for target acquisition and high frequency nodes for precise localization, recognition and tracking seems to be an advantage. This may include beam search instruction at higher frequencies from measurements at lower frequencies.
@@
Concerning the antennas and antenna arrays used, a wide variety of differences can occur. Here we only mention some. The difference between linear/planar vs. circular/cylindrical array, e.g., decides about full azimuth angle coverage which seems be more important for sensing than for communications. Moreover, elevation coverage matters if targets are expected in the air (e.g. drones). Important is also the difference between analog (phased array) and digital beamforming. While the former is capable only for beamforming, the latter has the potential of direction estimation. The reason is that estimation of direction is based on mutual correlation processing of array outputs. Analog beamforming (which seems dominating for higher frequencies) embodies a linear weight and sum processor that does not allow access to the individual antenna outputs for correlation processing. Full polarimetric processing requires polarization selective antennas and two corresponding UDC channels on the Tx and Rx side.
@@
The sniffer nodes discussed in section III may fall out of the ordinary even more. At first, a sniffer will act as receive only. But even if it is hosted by a gNB, it can cover extended frequency bands outside the host’s assigned bands. Receiving off-system signals would help to collect additional ISAC data in the reduced CPCL mode (like a synchronized passive radar) relative to other coexisting mobile radio networks that produce a valuable illuminating signal (could also include DVB-T broadcast). At the same time, the sniffer can be used to detect and localize alien emitters. This feature opens up other wide fields of ISAC applications and can be an interesting add-on for some use cases.
@@
If the sniffer is integrated into a distributed base station architecture by using a fixed wireless, RFoF, or optical fronthaul link, it can be directly incorporated into the internal gNB base-band processing regime and enjoy all advantages of CPCL processing. However, if it is equipped with its own subscriber identity (SIM card), it can act as normal UE and can be deployed more easily at preferred locations. Now it can interact with the hosting gNB via the normal UE UL/DL radio interface. Addionally, the CPCL advantages in terms of carrier and clock reference synchronization still apply.